that only 20 percent of the on 40 percent are paid for in the month after . The the month, I 300,000 330,000 450,00e Required: 3. Prepare a schedule that the firms financing cash flows for January through March
. The 31, 20x0 e sheet disclosed the following selected figures cash, $60.000; accounts receivable. $180.000: the month. I s paid at the time of repaying principal and computed on the portion of principal repaid at that time. Required: t Prepare a schedule that discloses the firms total cash collections for January through March. 3. Prepare a schedule that the firms financing cash flows for January through March References
media%2F4cd%2F4cd2f433-80e8-4fa0-943a-2d
. The 31, 20x0. at the the . Additional data month. Interest is paid at the time of repaying principal and computed on the portion of principal repaid at that time. 00,008 75,000 18,000 the firms
Answer #1
1) Cash collection:
Jan Feb Mar
coll of AR 36000 20% of AR
of jan sales 480000*0.6=288000 480000*0.3=144000
of feb sales 342000 171000
of mar sales 351000
sale of eqiup 18000
total cash collect 324000 486000 540000
2) Cash disbursements:
Jan Feb Mar
Payment of AP 69000
of Jan pur 300000*0.6=180000 120000
of feb pur 198000 132000
of mar pur 270000
Cash Ope costs' 96000 75000 138000
Total disbursements 345000 393000 540000
3) Cash Budget :
Jan Feb Mar
Open balance 60000 60000 131842.5
add: collections 324000 486000 540000
Total Cash available 384000 546000 671842.5
Less: Disbursements: 345000 393000 540000
Cash excess (def.) before financing 39000 153000 131842.5
Financing:
Borrowing 21000
less:Repayments 21000
less: Interest paid 157.5 21000*9%*1/12
Ending Cash Bal. 60000 131842.5 131842.5